how do you simplify linear equations like y=3/2x+2
Answer by Alan3354(69443)

how do you simplify linear equations like y=3/2x+2
--------------
That's as simple as that one gets. It's in slope-intercept form.
It can by put into "standard form"
y=3/2x+2
2y = 3x + 4
3x - 2y = -4 or
3x - 2y + 4 = 0
------------------
Standard form has the x and y terms on the same side, and uses integers.
"Simplified" can be a matter of opinion.